How long does it take to fill a swimming pool 9•5mx21•0m whose average depth is 2•5m of water that flows from 1•9cm diameter hose at a speed of 1•5m/s
1
Expert's answer
2017-12-13T12:05:07-0500
We are going to fill the volume V= 9.5m×21.0m×2.5m=498.75 m^3. The volume of water per second is the following: π×((1.9×〖10〗^(-2) m)/2)^2×1.5 m=4.25×〖10〗^(-4) m^3. It means that the total times will be t=(498.75m^3)/(4.25 m^3 s^(-1) )=13.6 days